a ball is thrown upward with an initial velocity of 20 m/s from the rooftop of a 16 meter tall building and falls freely until it hits the ground. what is the ball's final velocity just before hitting the ground


1
Expert's answer
2021-10-01T12:18:28-0400

Find the height the ball will reach above the roof:


"h=\frac{v^2}{2g}."

The velocity at the impact:


"u=\sqrt{2g(H+h)}=\sqrt{2g\bigg(H+\frac{v^2}{2g}\bigg)}=27\text{ m/s}."